OpenStudy (anonymous):

What is the probability of landing on heads on the 200 th flip on a coin?

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

1/2

OpenStudy (anonymous):

where is your work

Vocaloid (vocaloid):

probability of getting heads is always 1/2, assuming we are flipping a single coin.

Vocaloid (vocaloid):

one flip does not affect the probabilility of the second flip, or the third flip, or any other flip

OpenStudy (anonymous):

\[\frac{ C }{ 200 } So narrow \it down get \it alone.\]

Vocaloid (vocaloid):

now, if the question had asked, "what is the probability of getting 200 heads in a row" that question would be a bit different

OpenStudy (anonymous):

\[\frac{ C }{ 200}=\frac{ 1 }{ 2}\]

OpenStudy (anonymous):

\[\frac{ C }{ 200 } \]\[-\frac{ C }{ 100 }\] \[C=1\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Hope that helped

OpenStudy (anonymous):

=)